# Compression and complexity for sumset sizes in additive number theory

### v1

@@ Abstract.
The study of sums of finite sets of integers has mostly concentrated on sets with small sumsets (Freiman’s theorem and related work) and on sets with large sumsets (Sidon sets and $B_{h}$-sets). This paper considers the sets $\mathcal{R}_{\mathbf{Z}}(h,k)$ and $\mathcal{R}_{\mathbf{Z}^{n}}(h,k)$ of all sizes of $h$-fold sums of sets of $k$ integers or of $k$ lattice points, and the geometric and computational complexity of the sets $\mathcal{R}_{\mathbf{Z}}(h,k)$ and $\mathcal{R}_{\mathbf{Z}^{n}}(h,k)$.
@@
Sumset, restricted sumset, distribution of of sumset sizes, additive number theory, combinatorial number theory
@@ 1. Geometric and computational complexity
A central problem in additive number theory is to compute and understand the sumset size set $\mathcal{R}_{X}(h,k)$. This is unsolved even in the classical cases when $X$ is the semigroup $\mathbf{N}_{0}$ of nonnegative integers, the groups $\mathbf{Z}$, $\mathbf{Q}$, and $\mathbf{R}$ of integers, rational numbers, and real numbers, and the groups $\mathbf{Z}^{n}$, $\mathbf{Q}^{n}$, and $\mathbf{R}^{n}$ of lattice points with integral, rational, and real coordinates. For all positive integers $h$ and $k$, affine equivalence shows that
@@
We prove (Theorem 5) that
$\mathcal{R}_{\mathbf{Z}}(h,k)=\mathcal{R}_{\mathbf{Z}^{n}}(h,k).$
Also,
$\mathcal{R}_{\mathbf{N}_{0}}(h,k)\subseteq\mathcal{R}_{\mathbf{R}}(h,k)$
@@
$\min\mathcal{R}_{\mathbf{N}_{0}}(h,k)=\min\mathcal{R}_{\mathbf{R}}(h,k)=hk-h+1$
and
$\max\mathcal{R}_{\mathbf{N}_{0}}(h,k)=\max\mathcal{R}_{\mathbf{R}}(h,k)=\binom{h+k-1}{k-1}$
but it is not known if $\mathcal{R}_{\mathbf{N}_{0}}(h,k)=\mathcal{R}_{\mathbf{R}}(h,k)$.
For $u,v\in\mathbf{R}$, define the integer interval
$[u,v]=\{n\in\mathbf{Z}:u\leq n\leq v\}.$
@@
$\mathcal{R}_{\mathbf{Z}}(2,k)=\left[2k-1,\binom{k+1}{2}\right].$
@@
If sets $A$ an $A^{\prime}$ are affinely equivalent, that is, if
$A^{\prime}=\lambda\ast A+\mu=\{\lambda a+\mu:a\in A\}$
for real numbers $\lambda\neq 0$ and $\mu$, then $|hA^{\prime}|=|hA|$. This implies (1) and (2).
The diameter of a nonempty subset $A$ of a metric space is $\operatorname{\text{diam}}(A)=\sup\{d(x,y):x,y\in A\}$. Let $A=\{a_{1},a_{2},\ldots,a_{k-1},a_{k}\}$ be a set of $k$ nonnegative integers with $0=a_{1}<a_{2}<\cdots<a_{k-1}<a_{k}$ and $|hA|=t\in\mathcal{R}_{\mathbf{Z}}(h,k)$. If $a_{k}>ha_{k-1}+1$, then the “compressed” set $A^{\prime}=\{a_{1}a_{2},\ldots,a_{k-1},a^{\prime}_{k}\}$ with $a^{\prime}_{k}=ha_{k-1}+1$ also satisfies $|A^{\prime}|=k$ and $|hA^{\prime}|=t$ (Theorem 2), and
@@
For example, if $A=\{0,1,3,20,200\}$, then $|A|=5$ and $|4A|=65$. If $A^{\prime}=\{0,1,3,20,81\}$, then $|A^{\prime}|=5$ and $|4A^{\prime}|=65$. Indeed, for all $b\geq 81$, if $A^{\prime}=\{0,1,3,20,b\}$, then $|A^{\prime}|=5$ and $|4A^{\prime}|=65$. However, if $A^{\prime}=\{0,1,3,20,80\}$, then $|A^{\prime}|=5$ and $|4A^{\prime}|=80$.
@@
The set $\mathcal{R}_{\mathbf{Z}}(h,k)$ is finite. For all $t\in\mathcal{R}_{\mathbf{Z}}(h,k)$, there is an integer $N_{t}$ and a set $A\subseteq[0,N_{t}-1]$ with $|A|=k$ and $|hA|=t$. Let $N=\max\{N_{t}:t\in\mathcal{R}_{\mathbf{Z}}(h,k)\}$. Then
@@
$N(2,k)\leq 2^{k}.\operatorname{\qquad\text{and}\qquad}N(3,3)=5$
Problem 3.
@@ 2. Compression theorem for integers
(4) $a^{\prime}_{\ell}=\begin{cases}a_{\ell}&\text{if $\ell\in[1,j]$}\\ a_{\ell}-\delta_{j}&\text{if $\ell\in[j+1,k]$ }\end{cases}$
@@
The $h+1$ sets $(h-i)B+iC$ are pairwise disjoint if, for all $i\in[0,h-1]$,
@@
$(h-i-1)(a_{j}-a_{1})+i(a_{k}-a_{j+1})<a_{j+1}-a_{j}.$
This follows from inequality (2) and so the $h+1$ sets $(h-i)B+iC$ are pairwise disjoint and
@@
$a^{\prime}_{k}-a^{\prime}_{j+1}$ $=a_{k}-a_{j+1}$
@@
$a^{\prime}_{j+1}-a^{\prime}_{j}$ $=a_{j+1}-a_{j}-\delta_{j}$
@@
For all $i\in[0,h-1]$, the sets $(h-i)B^{\prime}+iC^{\prime}$ and $(h-1-i)B^{\prime}+(i+1)C^{\prime}$ are pairwise disjoint if
$(h-i)a_{j}+ia_{k}-i\delta_{j}<(h-1-i)a_{1}+(i+1)a_{j+1}-(i+1)\delta_{j}$
@@
If $a_{k}>1+ha_{k-1}$, then
$a_{k}-a_{k-1}>1+(h-1)a_{k-1}=1+(h-1)\max(a_{k-1}-1,a_{k}-a_{k})$
and the construction in Theorem 1 produces the set $A^{\prime}$ with
$a^{\prime}_{k}-a_{k-1}=1+(h-1)\max(a_{k-1}-a_{1},a_{k}-a_{k})=1+(h-1)a_{k-1}$
@@
If the set $A$ does not satisfy condition (3) for some $j\in[1,k-1]$, then apply Theorem 3 and obtain a set $A^{\prime}$ with $|hA^{\prime}|=|hA|$ and, by inequality (7), $\operatorname{\text{diam}}(A^{\prime})<\operatorname{\text{diam}}(A)$. Repeat this argument if the set $A^{\prime}$ does not satisfy inequality (10) for some $j\in[1,k-1]$. Continuing this procedure, we obtain a sequence of sets whose diameters form a strictly decreasing sequence of positive integers, and so the process must stop after finitely many iterations. This completes the proof. ∎
@@
for all $j\in[1,h-1]$, There is a unique integer $r$ in $[2,k-1]$ such that
@@ 3. Additive isomorphism $\mathcal{R}_{\mathbf{Z}^{n}}(h,k)=\mathcal{R}_{\mathbf{Z}}(h,k)$
If $f:A\rightarrow B$ is a Freiman isomorphism of order $h$, then $|hA|=|hB|$. For applications of Freiman isomorphisms, see [7].
@@
The proof of the following result is a standard application of Freiman isomorphisms.
@@
$g>n\max\{\|\mathbf{a}\|_{\infty}:\mathbf{a}\in A\}.$
@@ 4. Compression theorem for sums of sets of lattice points
Let $(\mathbf{x},\mathbf{y})$ be an inner product on $\mathbf{R}^{n}$. Let $\mathbf{u}$ be a unit vector in $\mathbf{R}^{n}$. For all $r\in\mathbf{R}$, there is the hyperplane
@@
$\geq(h-1)((\mathbf{u},\mathbf{a}_{k}-\mathbf{a}_{j_{1}+1}))$
@@
$\geq(h-1)((\mathbf{u},\mathbf{a}_{j_{2}}-\mathbf{a}_{1}))$
@@
Relation (14) implies that
@@
$(h-1-i)(r-q)+i(t-s)<s-r$
then, by (13), the sets $(h-i)B+iC$ and $(h-1-i)B+(i+1)C$ are disjoint.
@@
and so the sets $(h-i)B+iC$ are pairwise disjoint for all $i\in[0,h]$. It follows that
@@
For $\ell\in[j+1,k]$, we have
$s-\delta_{j}$ $=(\mathbf{u},\mathbf{a}_{j+1})-\delta_{j}=(\mathbf{u},\mathbf{a}_{j+1}-\delta_{j}\mathbf{u})=(\mathbf{u},\mathbf{a}^{\prime}_{j+1})$
@@
Because the standard basis vectors $\mathbf{e}_{1},\ldots,\mathbf{e}_{n}$ are pairwise orthogonal, the monotonicity inequalities (20) persist if each vector $\mathbf{a}_{j}$ is replaced by a vector of the form $\mathbf{a}_{j}-\mathbf{v}_{j}$ for any vector $\mathbf{v}_{j}$ in the hyperplane $H_{\mathbf{e}_{p}}(0)$, that is, by any vector that is a linear combination of the $n-1$ vectors in the set $\{\mathbf{e}_{1},\ldots,\mathbf{e}_{n}\}\setminus\{\mathbf{e}_{p}\}$.
@@ 5. An upper bound for $N(h,k)$
For all $\mathbf{x},\mathbf{y}\in(\mathbf{Z}/p\mathbf{Z})^{k}$, we have
$\|\mathbf{x}\pm\mathbf{y}\|_{\infty}\leq\|\mathbf{x}\|_{\infty}+\|\mathbf{y}\|_{\infty}.$
@@
From inequality (24), we have

### v2

@@ Abstract.
The study of sums of finite sets of integers has mostly concentrated on sets with small sumsets (Freiman’s theorem and related work) and on sets with large sumsets (Sidon sets and $B_{h}$-sets). This paper considers the sets ${\mathcal{R}}_{\mathbf{Z}}(h,k)$ and ${\mathcal{R}}_{{\mathbf{Z}}^{n}}(h,k)$ of all sizes of $h$-fold sums of sets of $k$ integers or of $k$ lattice points, and the geometric and computational complexity of the sets ${\mathcal{R}}_{\mathbf{Z}}(h,k)$ and ${\mathcal{R}}_{{\mathbf{Z}}^{n}}(h,k)$. For sumsets $hA$ with large diameter, there is a compression algorithm to construct sets $A^{\prime}$ with $|hA^{\prime}|=|hA|$ and small diameter.
@@
Sumset, restricted sumset, distribution of of sumset sizes, geometric compression, computational complexity, additive number theory, combinatorial number theory
@@ 1. Geometric and computational complexity
A central problem in additive number theory is to compute and understand the sumset size set $\mathcal{R}_{X}(h,k)$. This is unsolved even in the classical cases when $X$ is the semigroup $\mathbf{N}_{0}$ of nonnegative integers, the groups $\mathbf{Z}$, $\mathbf{Q}$, and $\mathbf{R}$ of integers, rational numbers, and real numbers, and the groups $\mathbf{Z}^{n}$, $\mathbf{Q}^{n}$, and $\mathbf{R}^{n}$ of lattice points with integral, rational, and real coordinates. If sets $A$ an $A^{\prime}$ are affinely equivalent, that is, if
$A^{\prime}=\lambda\ast A+\mu=\{\lambda a+\mu:a\in A\}$
for real numbers $\lambda\neq 0$ and $\mu$, then $|hA^{\prime}|=|hA|$. It follows that, for all positive integers $h$ and $k$,
@@
We have
(3) $\min\mathcal{R}_{\mathbf{Z}}(h,k)=\min\mathcal{R}_{\mathbf{R}}(h,k)=hk-h+1$
@@
(4) $\max\mathcal{R}_{\mathbf{Z}}(h,k)=\max\mathcal{R}_{\mathbf{R}}(h,k)=\binom{h+k-1}{k-1}$.
In Theorem 5, we prove that
$\mathcal{R}_{\mathbf{Z}}(h,k)=\mathcal{R}_{\mathbf{Z}^{n}}(h,k)$
for all positive integers $n$. Of course,
$\mathcal{R}_{\mathbf{Z}}(h,k)\subseteq\mathcal{R}_{\mathbf{R}}(h,k)$
but it is not known if $\mathcal{R}_{\mathbf{Z}}(h,k)=\mathcal{R}_{\mathbf{R}}(h,k)$ for some or all positive integers $h$ and $k$.
@@
$\mathcal{R}_{\mathbf{Z}}(2,k)=\left[2k-1,\binom{k+1}{2}\right$]
where, for all $u,v\in\mathbf{R}$,there is the integer interval
$u,v]=\{n\in\mathbf{Z}:u\leq n\leq v\}.$
@@
The diameter of a nonempty subset $A$ of a metric space is $\operatorname{\text{diam}}(A)=\sup\{d(x,y):x,y\in A\}$. Let $A=\{a_{1},a_{2},\ldots,a_{k-1},a_{k}\}$ be a set of $k$ nonnegative integers with $0=a_{1}<a_{2}<\cdots<a_{k-1}<a_{k}$ and $|hA|=t\in\mathcal{R}_{\mathbf{Z}}(h,k)$. If $a_{k}>ha_{k-1}+1$, then the “compressed” set $A^{\prime}=\{a_{1},a_{2},\ldots,a_{k-1},a^{\prime}_{k}\}$ with $a^{\prime}_{k}=ha_{k-1}+1$ also satisfies $|A^{\prime}|=k$ and $|hA^{\prime}|=t$ (Theorem 2), and
@@
For example, if $A=\{0,1,3,20,200\}$, then $|A|=5$ and $|4A|=65$. If $A^{\prime}=\{0,1,3,20,81\}$, then $|A^{\prime}|=5$ and $|4A^{\prime}|=65$. Indeed, for all $b\geq 81$, if $A^{\prime}=\{0,1,3,20,b\}$, then $|A^{\prime}|=5$ and $|4A^{\prime}|=65$. However, if $A^{\prime}=\{0,1,3,20,80\}$, then $|A^{\prime}|=5$ and $|4A^{\prime}|=64$.
@@
There are also computational complexity problems for sumsets. By (3) and (4), the set $\mathcal{R}_{\mathbf{Z}}(h,k)$ is finite. For all $t\in\mathcal{R}_{\mathbf{Z}}(h,k)$, there is an integer $N_{t}$ and a set $A\subseteq[0,N_{t}-1]$ with $|A|=k$ and $|hA|=t$. Let $N=\max\{N_{t}:t\in\mathcal{R}_{\mathbf{Z}}(h,k)\}$. Then
@@
$N(2,k)\leq 2^{k}\operatorname{\qquad\text{and}\qquad}N(3,3)=5.$
@@ 2. Compression theorem for integers
[…]
@@
The $h+1$ sets (hi)B+iC(h-i)B+iC are pairwise disjoint if, for all $i\in[0,h-1]$,
@@
$(h-1-i)(a_{j}-a_{1})+i(a_{k}-a_{j+1})<a_{j+1}-a_{j}.$
This is inequality (2) and so the $h+1$ sets (hi)B+iC(h-i)B+iC are pairwise disjoint and
@@
$a^{\prime}_{k}-a^{\prime}_{j+1}$ $=(a_{k}-\delta_{j})-(a_{j+1}-\delta_{j})=a_{k}-a_{j+1}$
@@
$a^{\prime}_{j+1}-a^{\prime}_{j}$ $=(a_{j+1}-\delta_{j})-a_{j}$
@@
For all $i\in[0,h-1]$, the sets $(h-i)B^{\prime}+iC^{\prime}$ and $(h-i-1)B^{\prime}+(i+1)C^{\prime}$ are pairwise disjoint if
$(h-i)a_{j}+ia_{k}-i\delta_{j}<(h-i-1)a_{1}+(i+1)a_{j+1}-(i+1)\delta_{j}$
@@
If $a_{k}>1+ha_{k-1}$, then there is a positive integer $\delta_{k-1}$ such that
$a_{k}-a_{k-1}$ $=1+\delta_{k-1}+(h-1)a_{k-1}$
$=1+\delta_{k-1}+(h-1)\max(a_{k-1}-1,a_{k}-a_{k}).$
The construction in Theorem 1 produces the set $A^{\prime}=\{a_{1}<\cdots<a_{k-1}<a^{\prime}_{k}\}$ with
$a^{\prime}_{k}=a_{k}-\delta_{k-1}=1+ha_{k}$
@@
If the set $A$ does not satisfy condition (5) for some $j\in[1,k-1]$, then apply Theorem 1 and obtain a set $A^{\prime}$ with $|hA^{\prime}|=|hA|$ and $\operatorname{\text{diam}}(A^{\prime})<\operatorname{\text{diam}}(A)$. Repeat this argument if the set $A^{\prime}$ does not satisfy inequality (12) for some $j\in[1,k-1]$. Continuing this procedure, we obtain a sequence of sets whose diameters form a strictly decreasing sequence of positive integers, and so the process must stop after finitely many iterations. This completes the proof. ∎
@@
for all $j\in[1,h-1]$, There is a unique integer $r$ in [2,k1][2,k-1] such that
@@ 3. Additive isomorphism $\mathcal{R}_{\mathbf{Z}^{n}}(h,k)=\mathcal{R}_{\mathbf{Z}}(h,k)$
Thus, $f$ induces a bijection from $hA$ to $hB$ and so $|hA|=|hB|$. For applications of Freiman isomorphisms, see [7].
@@
The proof of the following result is a simple application of Freiman isomorphisms.
@@
$g>h\max\{\|\mathbf{a}\|_{\infty}:\mathbf{a}\in A\}.$
@@ 4. Compression theorem for sums of sets of lattice points
Let $(\mathbf{x},\mathbf{y})$ be an inner product on $\mathbf{R}^{n}$ and let $\mathbf{u}$ be a unit vector in $\mathbf{R}^{n}$. For all $r\in\mathbf{R}$, we have the hyperplane
@@
$\geq(h-1)(\mathbf{u},\mathbf{a}_{k}-\mathbf{a}_{j_{1}+1})$
@@
$\geq(h-1)(\mathbf{u},\mathbf{a}_{j_{2}}-\mathbf{a}_{1})$
@@
The additive relation (16) implies that
@@
(h1i)(rq)+i(ts)<sr(h-1-i)(r-q)+i(t-s)<s-r
then, by (15), the sets (hi)B+iC(h-i)B+iC and (h1i)B+(i+1)C(h-1-i)B+(i+1)C are disjoint.
@@
and so the sets (hi)B+iC(h-i)B+iC are pairwise disjoint for all $i\in[0,h]$. It follows that
@@
Because $\mathbf{u}$ is a unit vector, we have $\delta_{j}=(\mathbf{u},\delta_{j}\mathbf{u})$. For all $\ell\in[j+1,k]$, we have
$s-\delta_{j}$ $=(\mathbf{u},\mathbf{a}_{j+1})-\delta_{j}=(\mathbf{u},\mathbf{a}_{j+1})-(\mathbf{u},\ \delta_{j}\mathbf{u})$
$=(\mathbf{u},\mathbf{a}_{j+1}-\delta_{j}\mathbf{u})=(\mathbf{u},\mathbf{a}^{\prime}_{j+1})$
@@
Because the standard basis vectors $\mathbf{e}_{1},\ldots,\mathbf{e}_{n}$ are pairwise orthogonal, the monotonicity inequalities (22) persist if each vector $\mathbf{a}_{j}$ is replaced by a vector of the form $\mathbf{a}_{j}-\mathbf{v}_{j}$ for any vector $\mathbf{v}_{j}$ in the hyperplane $H_{\mathbf{e}_{p}}(0)$, that is, by any vector that is a linear combination of the n1n-1 vectors in the set $\{\mathbf{e}_{1},\ldots,\mathbf{e}_{n}\}\setminus\{\mathbf{e}_{p}\}$.
@@ 5. An upper bound for $N(h,k)$
The following result is known. A proof is included for completeness.
@@
For all $\mathbf{v}_{1},\mathbf{v}_{2}\in(\mathbf{Z}/p\mathbf{Z})^{k}$, we have
(23) $\|\mathbf{v}_{1}\pm\mathbf{v}_{2}\|_{\infty}\leq\|\mathbf{v}_{1}\|_{\infty}+\|\mathbf{v}_{2}\|_{\infty}.$
@@
From inequalities (23) and  (27), we have